OpenStudy (anonymous):

(x+2)^2 +11(x+2)+24 need solution set!!

myininaya (myininaya):

is this equal to zero?

myininaya (myininaya):

let u=x+2 so \[u^2=(u+2)^2\] u^2+11u+24=0 (u+8)(u+3)=0 \[u=-8,u=-3\] \[x+2=-8,x+2=-3\] \[x=-8-2,x=-3-2\] \[x=-10,x=-5\]

myininaya (myininaya):

the solution set is {-10,-5}

OpenStudy (anonymous):

Thank YOU

myininaya (myininaya):

:)

OpenStudy (anonymous):

0.2t=0.4+0.7t what is soulution set?

OpenStudy (anonymous):

I hope you can help with this one please

myininaya (myininaya):

0.2t-0.7t=0.4 -0.5t=0.4 t=0.4/(-0.5) t=-4/5